OpenStudy (anonymous):

Fine, since the runner doesn't have a fair play, let me do this EQUATION IS 250 x 588 / 8 x 0 The person with the correct answer gets a 3 fans and 12 medals (i have more than one account) See how you like them apples

OpenStudy (anonymous):

0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

1

OpenStudy (anonymous):

jk undefined

OpenStudy (anonymous):

Its division by zero

OpenStudy (ashleyisakitty):

What a simple equation for such a large prize

OpenStudy (anonymous):

0/1=0 1/0= undefined

OpenStudy (anonymous):

C MON ppl easy stuff!!!!

OpenStudy (anonymous):

I aid the first person didn't I? Last time I checked, I was a person! So, y'all were a good try. lol

OpenStudy (anonymous):

I deserve all the medals >.>

OpenStudy (anonymous):

lol. Well you didn't get it

OpenStudy (anonymous):

0/1=0 1/0= undefined yes i got it XD